OpenStudy (lilymq):

What is the solution of log 3x + 4^125 = 3?

OpenStudy (lilymq):

Is it 4?

OpenStudy (anonymous):

Remember the Log rule, log a (m) = x is the same as a^x = m In this question, a is 3x + 4, m is 125, x is 3. log 3x + 4 (125) = 3 3x + 4 ^ 3 = 125 3x + 4 ^ 3 = 5 ^3 3x + 4 = 5 3x = 1 x = 1/3

OpenStudy (lilymq):

thank you so much!

OpenStudy (anonymous):

your welcome :)